# Information About Other Agents in Mechanism Design

### v3

@@ Abstract
We study mechanism design settings where the planner has an interest in agents receiving noisy signals about the types of other agents. When the planner is interested only in partial implementation — designing a mechanism for which there exists an equilibrium that yields a desired outcome — any social choice rule that is incentive-compatible after agents receive additional information about other agents is incentive-compatible without this information. However, additional information about other agents can eliminate undesired equilibria, making it helpful to a planner interested in full implementation — designing a mechanism for which every equilibrium outcome is desirable. We provide a sufficient condition under which a social choice function that is not fully implementable when agents have no information about types of other agents can become fully implementable if agents have additional information.
@@ 1 Introduction
The goal of a social planner in any mechanism design problem is to design a game under which a desirable outcome emerges in equilibrium. Agents have private information about their own preferences and preferences of other agents. Naturally, the set of equilibrium outcomes that can emerge from a mechanism is sensitive to the details of the information available to agents. A considerable body of mechanism design literature focuses on the effects of changing information structures on achieving desirable outcomes. Most of this work focuses on how much agents know about variables that directly affect their own payoffs — for example, the information a bidder knows about her own valuation for an object (Bergemann et al. (2022)) — rather than information about other agents’ preferences.
In this paper, we examine how the planner can have an interest in agents possessing additional information about the preferences of other agents, revisiting the Bayesian implementation problem as studied by Jackson (1991) (see also Postlewaite and Schmeidler (1986) and Palfrey and Srivastava (1989)). In this setting, agents fully understand their own preferences over a set of outcomes but do not know the preferences of other agents. We consider how the implementability of a social choice function (SCF) changes if agents can receive a full-support signal regarding the preferences of other agents. In particular, our main result gives a sufficient condition under which a planner interested in full implementation — designing a mechanism such that every equilibrium yields a desirable outcome — will strictly prefer an agent to have additional information about agents.
In section 3, we show that any SCF which satisfies incentive compatibility when agents receive informative signals about the types of other agents also satisfies incentive compatibility when agents do not have this additional information. This implies that a planner concerned only with partial implementation — ensuring that the desirable outcome constitutes an equilibrium of the mechanism — will never strictly prefer an agent to have additional information about other agents. The argument is straightforward: When an agent has access to an informative signal about the preferences of other agents, the planner must ensure that the agent is incentivized to report truthfully in each signal realization. A weighted average of these constraints across signal realizations yields the incentive constraint needed for the agent if she did not have access to additional information, making this latter constraint easier to satisfy.
@@
The remainder of the paper examines how additional information about the types of other agents can help an agent signal untruthful equilibria to the planner. In section 4, we present the main example of the paper: an SCF that is not fully implementable when agents have no information about the preferences of other agents but is fully implementable when agents have more information. The condition that rules out undesired equilibria in Bayesian implementation settings is Bayesian monotonicity, which guarantees that agents have sufficient incentives to signal to the planner that an undesirable equilibrium strategy is being played. We focus heavily on how Bayesian monotonicity varies with the information agents have about other agents. Section 5 presents a sufficient condition for an SCF that is not fully implementable when agents have no information about the types of other agents to be fully implementable when agents have more information.
@@
To establish our main result, we note that an incentive-compatible SCF is not fully implementable if and only if there is an untruthful equilibrium of the direct mechanism such that no agent has an incentive to report the deception to the planner. The task is then to construct a signal that gives some agent $i$ the ability to eliminate these equilibria. In general, if truth-telling is a dominant strategy for player $i$, an untruthful equilibrium $\alpha$ of the direct mechanism will involve some subset of other agents playing deceptions. The signal we construct partitions the set of other agents’ types into two groups. Each group in the partition must contain a type profile of other agents such that if the agent $i$ knew this type profile realized, she would prefer all other agents to report truthfully rather than deceive according to $\alpha$, regardless of her own type. If agent $i$ is confident enough that the realized type profile is in one group or the other, she can propose an alternate SCF that swaps the outcomes implemented under the truthful strategy versus the deception. By construction, this alternate SCF can only give agent $i$ a weakly lower payoff relative to the original if all other agents report truthfully, so agent $i$ would only propose it if $\alpha$ is being played. Whenever the signal is sufficiently accurate, agent $i$ can eliminate $\alpha$ as an equilibrium. We conclude by showing that this procedure can iteratively eliminate all untruthful equilibria of the direct mechanism.
Though a planner may want agents to have access to informative signals about the types of other agents compared to no information, our analysis also demonstrates that a planner may prefer an information structure that lies strictly between no information and complete information. This conclusion is consistent with the rapidly growing literature on information design spurred by Rayo and Segal (2010) and Kamenica and Gentzkow (2011).
@@
This paper directly contributes to the literature on full Bayesian implementation — which began with Postlewaite and Schmeidler (1986), Palfrey and Srivastava (1989), and Jackson (1991) — by shedding light on the relationship between Bayesian monotonicity and the information available to agents. Since agents are receiving informative signals about the state of the world, we also add to the extensive literature on the effects of changing information structures in mechanism design settings pioneered by Bergemann and Välimäki (2002). It relates especially to the literature on information disclosure in mechanism design. In auction settings, Ganuza (2004) studies the optimal way for a seller to disclose a costly public signal about a good’s characteristic in a second-price auction, which is a component of bidders’ private valuations, while Bergemann and Pesendorfer (2007) investigate the joint design problem in which the seller controls the information bidders have about their own valuations and the auction rules. Other papers that have investigated the effects of information disclosure in auction settings include Eső and Szentes (2007), and Bergemann et al. (2022). Similar questions have also been investigated in screening settings by Li and Shi (2017) and Roesler and Szentes (2017). Since much of the existing literature studies cases when agents’ types are independently drawn, existing results focus on how much agents know about their own preferences. This paper focuses on a mechanism designer’s interest in agents’ information about types of other agents. Additionally, unlike the auction papers that portray the auctioneer as having the ability to provide information to agents, this paper remains agnostic on where the information about other agents comes from. The emphasis is on how changing information structures can affect the planner’s implementation goals. Finally, while our paper asks whether a planner can be interested in agents’ information about other agents, one can also ask about an agent’s incentives to acquire such information, as done in Gleyze and Pernoud (2023), who show that most mechanisms incentivize agents to acquire information about other agents in environments where agents have uncertainty about their own types as well as types of other agents.
@@ 2 Model
[…]
The planner wishes to implement a social choice function (SCF) $F:\Theta\to A.$ The planner commits to a mechanism $\Gamma=\langle M,h\rangle$, such that $M=\prod_{i=1}^{n}M_{i}$ consists of a message space $M_{i}$ for each agent $i$ and $h:M\to A$ is the outcome function. A strategy profile $s=(s_{1},\dots,s_{n})$ specifies for each agent $i$ a function $s_{i}:\Theta_{i}\to\Delta(M_{i})$. At any state $\theta$, let $s(\theta)=(s_{1}(\theta_{1}),\dots,s(\theta_{n}))$ be the message profile sent by all agents, and $s_{-i}(\theta_{-i})=(s_{j}(\theta_{j}))_{j\neq i}$ denote the messages sent by agents other than $i.$ Let $B^{\Gamma}$ be the set of Bayesian Nash equilibria for a mechanism $\Gamma.$ If there is a $s\in B^{\Gamma}$ such that $h\circ s=F$ then $\Gamma$ partially implements $F$, and $F$ is considered partially implementable. If $\Gamma$ partially implements $F$, and $B^{\Gamma}$ contains only one equilibrium outcome, then $\Gamma$ fully implements $F$ and $F$ is considered fully implementable.
Classical mechanism design results pin down two key conditions for the implementability of an SCF $F$ in Bayesian settings.
@@
A deception for agent $i$ is a function $\alpha_{i}:\Theta_{i}\to\Theta_{i}.$ A deception $\alpha=(\alpha_{1},\dots,\alpha_{n})$ consists of a deception $\alpha_{i}$ for each agent $i\in N.$
@@
It was shown in Jackson (1991) that incentive compatibility and Bayesian monotonicity are necessary and sufficient for an SCF $F$ to be fully implementable in economic environments with at least 3 agents. [^2] An environment is economic if given any state $\theta\in\Theta$ and SCF $F$, at least two agents prefer some outcome $a\in A$ to $F(\theta).$ [^3] Jackson (1991) also identifies a closure condition for collections of SCFs, social choice sets, but the condition is vacuous for the case of a single-valued SCF. While incentive compatibility guarantees that truth-telling is an equilibrium of the direct mechanism, Bayesian monotonicity is the condition that allows the planner to design a mechanism that can eliminate undesired equilibria. The intuition behind the condition, which we will reference throughout this paper, is as follows. Suppose agents are initially at an equilibrium that uses a deception $\alpha$ such that $F\circ\alpha\neq F.$ By allowing any individual agent $i$ to specify an alternate SCF (the function $y$ can be extended to $\Theta$ by implementing the same outcome regardless of agent $i$’s type) that is implemented only if it appears to harm agent $i$, an agent can choose to use this alternate SCF to signal to the planner that the agents are lying. Upon seeing the alternate SCF, the planner understands that a deception is being played because it would harm agent $i$ in a truth-telling equilibrium, regardless of her own type. Bayesian monotonicity gives at least one type of agent $i$ an incentive to specify an alternate SCF because she would prefer the alternate SCF to $F$ under the deception $\alpha$, ruling out the deception as an equilibrium. Bayesian monotonicity guarantees that such a procedure can be done for every undesirable deception.
@@
The timing of the game is as follows.
Nature chooses $\theta$ from $\mu$, planner announces $\Gamma$, Structure of signals$\{\sigma_{i}\}_{i\in I}$ announcedAgents observe realizations of $\sigma_{i}$gents choose strategies $s_{i}$gents learn $\theta_{i}$,send messagesaccording to $s_{i}$utcome realized
Figure 1: Mechanism timing.
@@
When the signals $\{\sigma_{i}\}_{i\in I}$ provide no information, the model is equivalent to the classical Bayesian implementation problem studied by Jackson (1991). The main question we address is whether and when an SCF $F$ that is not partially (resp. fully) implementable when $\sigma_{i}$ is noninformative for all $i$ can become partially (resp. fully) if one or more of the $\sigma_{i}$ is nontrivial. We are agnostic on where the signals $\{\sigma_{i}\}_{i\in I}$ come from. The nature of the question is one of comparative statics on the implementability of an SCF $F$ with respect to an increase in information one more agents $i$ has regarding $\Theta_{-i}$. [^4] Since agents initially do not have any information about $\Theta_{-i}$ aside from the common prior $\mu$, any nontrivial signal $\sigma_{i}$ represents an increase in agent $i$’s information with respect to the Blackwell order. In the remainder of this paper we will say that $F$ is partially implementable with respect to $\{\sigma_{i}\}_{i\in I}$ if truth-telling is an equilibrium in the direct mechanism in each realization of the signals $\{\sigma_{i}\}_{i\in I}$. Similarly define fully implementable with respect to $\{\sigma_{i}\}_{i\in I}.$ Note that when agents receive a signal $\sigma_{i}$ about other agents, both $\theta_{i}$ and $\sigma_{i}$ are both part of agent $i$’s private information, so if $\sigma_{i}$ has possible realizations in $\zeta_{i}=\{A_{i1},\dots,A_{ik}\}$ an implementing mechanism should ask each agent $i$ to specify a strategy for each $(\theta_{i},A_{ij})$ pair. Our setup is implicitly analyzing an equivalent formulation of this mechanism in which the uncertainty coming from the signals $\{\sigma_{i}\}$ is resolved first and asking whether $F$ is partially or fully implementable in each “interim game” of the full mechanism.
@@
Our setup makes several substantive assumptions. Other than requiring all $\sigma_{i}$ to have full support, we assume there are no restrictions over the structure of the signal agent $i$ can receive. Thus, agents do not have private information that can constrain the types of signals they can process. Additionally, we assume the signals are not costly in order to focus on the role of the additional information regarding other agents’ types. Finally, the structure of the signals $\{\sigma_{i}\}_{i\in I}$ is assumed to be common knowledge.
2.1 The full-support signal assumption
The restriction that each $\sigma_{i}$ has full support is a simplifying assumption that avoid complications to the type space $\Theta$ after agents receive their signals. To illustrate we make this assumption, consider the following example under which agents receive perfectly informative signals. Suppose the environment consists of three agents $i=1,2,3$, where each agent $i$ has types $\theta_{i},\theta_{i}^{\prime}.$ There is a uniform common prior, so all states are equally likely. There are 5 possible outcomes $A=\{a,b,c,d,e\}.$ The SCF $F$ is as follows.
$F\mid\theta_{3}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $a$
$\theta_{1}^{\prime}$ $a$ $b$
$F\mid\theta_{3}^{\prime}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $c$ $c$
$\theta_{1}^{\prime}$ $c$ $b$
This notation specifies $F$ by separating the types of agent 1 along rows, agent 2’s types along columns, and agent 3’s types along the left versus right boxes. For example, $F(\theta_{1},\theta_{2},\theta_{3})=a$ while $F(\theta_{1},\theta_{2},\theta_{3}^{\prime})=c.$ The agents have the following preferences.
Type Preferences
$\theta_{1}$ $a\succ b\succ c\succ d\succ e$
$\theta_{1}^{\prime}$ $d\succ c\succ b\succ a\succ e$
$\theta_{2}$ $c\succ a\succ b\succ d\succ e$
$\theta_{2}^{\prime}$ $b\succ a\succ c\succ d\succ e$
$\theta_{3}$ $e\succ a\succ b\succ c\succ d$
$\theta_{3}^{\prime}$ $e\succ c\succ b\succ a\succ d$
Further, suppose agent 1’s vNM utility function $u_{1}$ is such that:
$u_{1}(a,\theta_{1})=3,u_{1}(b,\theta_{1})=2,u_{1}(c,\theta_{1})=0,$
and
$u_{1}(c,\theta_{1}^{\prime})=3,u_{1}(b,\theta_{1}^{\prime})=1,u_{1}(a,\theta_{1}^{\prime})=0.$
Then $F$ does not satisfy incentive compatibility in this information structure, since
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=\frac{3}{2}<\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1})\mid\theta_{1})=\frac{7}{4},$
and
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=\frac{5}{4}<\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=\frac{3}{2}.$
If $\sigma_{i}$ is uninformative for all $i$, it follows that $F$ is not fully or partially implementable under the specified information structure. On the other hand, if $\sigma_{i}$ is perfectly informative for all $i$, each agent $i$ has complete information regarding the state $\theta$ upon observing the realization of $\sigma_{i}$ and learning their own type $\theta_{i}$. Then incentive compatibility becomes trivial when $n\geq 3$. Each agent’s private information now consists of the complete state $\theta$. Thus if one agent unilaterally deviates from reporting their private information truthfully, the planner can simply ignore the deviation and implement the outcome in the state reported by $n-1$ agents. In this case, full implementability of $F$ reduces to the classical Nash implementation problem solved by Maskin (1999), which proves that Maskin monotonicity [^5] An SCF $F$ is Maskin-monotonic if and only if for all $\theta,\theta^{\prime}\in\Theta$, whenever $F(\theta)\neq F(\theta^{\prime})$, there exist $i\in I$ and $a\in A$ such that $u_{i}(F(\theta),\theta)\geq u_{i}(a,\theta)$ but $u_{i}(F(\theta),\theta^{\prime})<u_{i}(a,\theta^{\prime})$. and no veto power [^6] An SCF $F$ satisfies no veto power if for all $(a,j,\theta)\in A\times I\times\Theta$, the condition $u_{i}(a,\theta)\geq u_{i}(b,\theta)$ for all $b\in A$ and $i\neq j$ implies $a=F(\theta).$ are sufficient conditions for the full implementability of $F$ when there are at least three agents. No veto power is automatically satisfied because $F$ never allows two agents to simultaneously get their most preferred outcome. To see that $F$ is Maskin-monotonic, first note that outcomes $a$ and $c$ fall relative to each other in agent 3’s preference ranking when moving between types $\theta_{3}$ and $\theta_{3}^{\prime}$ and vice versa. Additionally, $b$ falls relative to $a$ in agent 1’s preference ranking when moving from type $\theta_{1}^{\prime}$ to $\theta_{1}$, and $a$ and $c$ both fall relative to $d$ when agent 1 moves from type $\theta_{1}$ to $\theta_{1}^{\prime}.$ Finally, $a$ and $c$ fall relative to each other in agent 2’s preference ranking when moving between types $\theta_{2}$ and $\theta_{2}^{\prime}$ and vice versa, and $b$ falls relative to $a$ and $c$ when moving from $\theta_{2}^{\prime}$ to $\theta_{2}.$ Thus $F$ is implementable if each $\sigma_{i}$ provides agent $i$ with a perfect signal regarding $\Theta_{-i}.$
This example demonstrates that agents receiving information about the preferences of others can allow an initially non-implementable SCF to become implementable. It also shows how the incentive compatibility condition changes when agents are able to completely rule out certain states of the world. In particular, since implementation with complete information is a special case of Bayesian implementation, it must be that $F$ satisfies incentive compatibility under complete information even though it is not incentive-compatible under the uniform common prior. This is because an agent’s type $\theta_{i}$ encapsulates information about both the preferences of agent $i$ and the information she has about the types of other agents. In the example above, $\theta_{i}$ is a complete description of agent $i$’s type under the uniform common prior, because it specifies her preferences over the outcomes but she is not able to rule out any type profile of the other agents. Under complete information, however, $\theta_{i}$ is no longer an adequate description of her type. Since agent $i$ is able to completely discern the types of the other agents, her type includes what she knows to be the type profile of the other agents. As such, if one were to translate the example above with complete information into the Bayesian implementation setup, then $\Theta_{i}=\prod_{j=1}^{3}\{\theta_{j},\theta_{j}^{\prime}\}$ for all $i.$ So, when agent $i$ has preferences according to $\theta_{i},$ she has four corresponding types $t_{i}\in\Theta_{i}$. This is consistent with the mechanism exhibited in Maskin (1999), where each agent was asked to report the complete state of the world, which included information about the preferences of all agents.
In the translation of a complete information implementation problem into one of Bayesian implementation, the states $\Theta=\prod_{i}\Theta_{i}$ contain many zero-probability states. In fact, any $t=(t_{1},t_{2},t_{3})\in\Theta$ for which each $t_{1}=t_{2}=t_{3}$ does not hold is a state that occurs with zero probability. However, the planner still specifies the SCF for these events in order to preclude equilibria that are deceptions, which trivializes incentive-compatibility when $n\geq 3$. As such, the restriction that each $\sigma_{i}$ always has full support simply prevents changes to any type space $\Theta_{i}$ after agents receive the signal realization.

### v4

@@ Abstract
We study mechanism design settings where the planner has an interest in agents receiving noisy signals about the types of other agents. We show that additional information about other agents can eliminate undesired equilibria, making it helpful to a planner interested in full implementation — designing a mechanism for which every equilibrium outcome is desirable. We provide a sufficient condition under which a social choice function that is not fully implementable when agents have no information about types of other agents can become fully implementable when agents have additional information.
@@ 1 Introduction
A key challenge of mechanism design is to design a game under which privately informed agents reach a desired outcome in equilibrium. The effectiveness of these mechanisms often depends critically on the information available to agents. Many recent papers have studied how varying what agents know about variables that directly affect their own payoffs — for example, the information a bidder knows about her own valuation for an object (Bergemann et al. (2022)) — affects a designer’s goals. But in many settings agents naturally possess or can acquire information about other agents as well. Can a social planner strictly benefit from agents having additional information about other agents in the mechanism? We address this question within the canonical framework of Bayesian implementation of Jackson (1991) (see also Postlewaite and Schmeidler (1986) and Palfrey and Srivastava (1989)). Our main finding is that a planner interested in full implementation — designing a mechanism such that every equilibrium yields a desirable outcome — can strictly prefer some agents to receive noisy signals about other agents, and we give a sufficient condition under which this can occur.
As an example, consider a firm deciding whether to invest more in a product, maintain the status quo, or exit the market. The board (planner) relies on two analysts who privately observe positively correlated signals about whether demand is high or low. One analyst is more senior than the other, and the board is cautious: they want the firm to exit whenever the senior analyst observes a low signal, invest only when both analysts observe a high signal, and maintain the status quo if the senior analyst believes the demand is high and the junior analyst disagrees. Assuming the analysts care only about monetary payments, a natural way to ensure both analysts’ truthful reporting of their signal as an equilibrium is to pay the analysts if and only if they agree. Unfortunately, this setup admits undesirable equilibria, such as analysts coordinating on a high or low demand report, regardless of their private signals. A key theoretical insight from the Bayesian implementation literature is that allowing agents to suggest alternative outcome rules can eliminate undesired equilibria if the agent proposing the alternate rule strictly benefits from the proposal at the undesired equilibrium but is weakly worse off if agents are at the preferred equilibrium. To illustrate our main point, we introduce two more agents — a CEO and a compliance officer — who do not have private information but have preferences over the firm’s action. We exhibit a common prior and natural payoff structure under which the compliance officer can propose alternate rules that eliminate the equilibria where the analysts always coordinate on a message, but no agent is able to rule out an equilibrium where the analysts always report the demand is high when they observe a low signal and vice versa.
The CEO wants the firm to capture as large a share of the product’s market as possible, so she prefers investing to the status quo, and exiting is her least-preferred outcome. Suppose we give the CEO a binary signal about the analysts’ private information that achieves the following. In one realization, she assigns higher probability to both analysts seeing a high signal than to both seeing a low signal. In the other realization, she assigns higher probability to the senior analyst observing a high signal and the junior analyst observing a low signal than vice versa. In either realization, she can credibly propose an alternative outcome rule that swaps the firm actions prescribed to the respective pair of analyst signals. Each swap weakens her expected payoff when the analysts are truthful, so she never invokes it in the desired equilibrium. But it strictly improves her payoff when the analysts are engaged in the deceptive equilibrium where they always lie, giving her an incentive to propose the alternate rule and thereby break that equilibrium.
@@
Motivated by this example, this paper seeks conditions under which a planner would strictly prefer that agents have full-support signals about other agents’ types. While the above example provides intuition for why such information can help, our starting result in Section 3 is that additional information never strictly benefits a planner interested only in partial implementation — merely ensuring truth-telling is an equilibrium. The remainder of the paper examines how additional information about the types of other agents can help an agent signal untruthful equilibria to the planner. In section 4, we formally describe the example from above. We show that the board’s desired social choice function (SCF) is not fully implementable when agents have no information about the types of other agents but is fully implementable when the CEO is given a noisy signal. The condition that rules out undesired equilibria in Bayesian implementation settings is Bayesian monotonicity, which guarantees agents sufficient incentives to signal undesirable equilibria to the planner. Our theoretical analysis focuses heavily on how Bayesian monotonicity varies with the information agents have about other agents. Section 5 presents a sufficient condition for an SCF that is not fully implementable when agents have no information about the types of other agents to be fully implementable when agents have more information.
@@
To establish our main result, we note that a partially implementable SCF is not fully implementable if and only if there is an untruthful equilibrium of the direct mechanism such that no agent has an incentive to report the deception to the planner. The task is then to construct a signal that gives some agent $i$ the ability to eliminate these equilibria. In general, if truth-telling is a dominant strategy for player $i$, an untruthful equilibrium $\alpha$ of the direct mechanism will involve some subset of other agents playing deceptions. The signal we construct partitions the set of other agents’ types into two groups. Each group in the partition must contain a type profile of other agents such that if the agent $i$ knew this type profile realized, she would prefer all other agents to report truthfully rather than deceive according to $\alpha$, regardless of her own type. If agent $i$ is confident enough that the realized type profile is in one group or the other, she can propose an alternate SCF that swaps the outcomes implemented under the truthful strategy versus the deception. By construction, this alternate SCF can only give agent $i$ a weakly lower payoff relative to the original if all other agents report truthfully, so agent $i$ would only propose it if $\alpha$ is being played. Whenever the signal is sufficiently accurate, agent $i$ can eliminate $\alpha$ as an equilibrium. We conclude by showing that this procedure can iteratively eliminate all untruthful equilibria of the direct mechanism.
@@
This paper directly contributes to the literature on full Bayesian implementation — which began with Postlewaite and Schmeidler (1986), Palfrey and Srivastava (1989), and Jackson (1991) — by shedding light on the relationship between Bayesian monotonicity and the information available to agents. Since agents are receiving informative signals about the state of the world, we also add to the extensive literature on the effects of changing information structures in mechanism design settings pioneered by Bergemann and Välimäki (2002). It relates especially to the literature on information disclosure in mechanism design. In auction settings, Ganuza (2004) studies the optimal way for a seller to disclose a costly public signal about a good’s characteristic in a second-price auction, which is a component of bidders’ private valuations, while Bergemann and Pesendorfer (2007) investigate the joint design problem in which the seller controls the information bidders have about their own valuations and the auction rules. Other papers that have investigated the effects of information disclosure in auction settings include Eső and Szentes (2007), and Bergemann et al. (2022). Similar questions have also been investigated in screening settings by Li and Shi (2017) and Roesler and Szentes (2017). Additionally, while our paper asks whether a planner can be interested in agents’ information about other agents, one can also ask about an agent’s incentives to acquire such information, as done in Gleyze and Pernoud (2023), who show that most mechanisms incentivize agents to acquire information about other agents in environments where agents have uncertainty about their own types as well as types of other agents. Finally, our emphasis on a designer’s preferences over agent information structures connects with the recent literature on information design spurred by Rayo and Segal (2010) and Kamenica and Gentzkow (2011). Consistent with the information design literature, our analysis shows that a designer may prefer information structures strictly between no information and complete information.
@@ 2 Model
There are $n\geq 3$ agents $I=\{1,\dots,n\}$, a finite set of outcomes $A$, and a finite number of payoff types (henceforth types) $\theta_{i}\in\Theta_{i}$ for each agent $i.$ The state of the world $\theta\equiv(\theta_{1},\dots,\theta_{n})$ is in the set $\Theta\equiv\Theta_{1}\times\cdots\times\Theta_{n}.$ For any agent $i$, the type profile of agents other than $i$ is denoted $\theta_{-i}\in\Theta_{-i}\equiv\prod_{j\neq i}\Theta_{j}.$ There is a full-support common prior $\mu$ over $\Theta$, and each agent knows her own type $\theta_{i}$ but cannot fully determine $\theta_{-i}.$ Each agent $i\in I$ has a utility function that depends only on $\Theta_{i}$, so $u_{i}:A\times\Theta_{i}\to{\mathbb{R}}$. Thus, for every type $\theta_{i}\in\Theta_{i}$, agent $i$ has a set of preferences over the outcomes in $A.$ This is the case of private values.
The planner wishes to implement a deterministic social choice function (SCF) $F:\Theta\to A.$ The planner commits to a mechanism $\Gamma=\langle M,h\rangle$, such that $M=\prod_{i=1}^{n}M_{i}$ consists of a message space $M_{i}$ for each agent $i$ and $h:M\to A$ is the outcome function. A strategy profile $s=(s_{1},\dots,s_{n})$ specifies for each agent $i$ a function $s_{i}:\Theta_{i}\to M_{i}$. [^2] The restriction to pure strategies is only for expositional simplicity and is discussed in footnote 5. At any state $\theta$, let $s(\theta)=(s_{1}(\theta_{1}),\dots,s(\theta_{n}))$ be the message profile sent by all agents, and $s_{-i}(\theta_{-i})=(s_{j}(\theta_{j}))_{j\neq i}$ denote the messages sent by agents other than $i.$ Let $B^{\Gamma}$ be the set of Bayesian Nash equilibria for a mechanism $\Gamma.$ If there is an $s\in B^{\Gamma}$ such that $h\circ s=F$ then $\Gamma$ partially implements $F$, and $F$ is considered partially implementable. If $\Gamma$ partially implements $F$, and $B^{\Gamma}$ contains only one equilibrium outcome, then $\Gamma$ fully implements $F$ and $F$ is considered fully implementable. [^3] Standard implementation theory allows the planner to specify a social choice set (SCS). Full implementation of an SCS is equivalent to full implementability of each SCF in the set under a full-support prior. We assume the environment is economic: For any state $\theta\in\Theta$ and any SCF $F$, at least two agents would prefer to alter the outcome prescribed by $F.$
Classical mechanism design results pin down two key conditions for the implementability of an SCF $F$.
@@
A deception for agent $i$ is a function $\alpha_{i}:\Theta_{i}\to\Theta_{i}.$ A deception $\alpha=(\alpha_{1},\dots,\alpha_{n})$ consists of a deception $\alpha_{i}$ for each agent $i\in I.$
@@
It was shown in Jackson (1991) that incentive compatibility and Bayesian monotonicity are necessary and sufficient for an SCF $F$ to be fully implementable in economic environments with at least 3 agents. [^4] Jackson (1991) identifies a closure condition necessary for the implementation of SCSs, but the condition is vacuous for the case of a single-valued SCF. [^5] Jackson (1991) characterizes full implementation in pure strategies, i.e. every pure-strategy equilibrium of the mechanism is desirable. If agents can use mixed strategies, full implementation in our setting should demand that every realization of the strategy profile yields an outcome that aligns with $F.$ Serrano and Vohra (2010) identify a mixed Bayesian monotonicity condition that together with incentive compatibility is necessary and sufficient for full implementation of an SCF in economic environments when agents can use mixed strategies. The condition is equivalent to Bayesian monotonicity in the case of a single-valued SCF, implying that all our results hold if we allow agents to use mixed strategies. While incentive compatibility guarantees that truth-telling is an equilibrium of the direct mechanism, Bayesian monotonicity is the condition that allows the planner to design a mechanism that can eliminate undesired equilibria. The intuition behind the condition, which we will reference throughout this paper, is the following. Suppose agents are initially at an equilibrium that uses a deception $\alpha$ such that $F\circ\alpha\neq F.$ The planner can enlarge the message space of each agent beyond her own type space to allow any individual agent $i$ to specify an alternate SCF. (The function $y$ in the definition of Bayesian monotonicity can be extended to $\Theta$ by implementing the same outcome regardless of agent $i$’s type.) The planner can commit to implementing the alternate SCF only if it appears to weakly harm agent $i$, regardless of her type, if the agents were playing the truthful equilibrium. Bayesian monotonicity gives at least one type of agent $i$ an incentive to specify such an alternate SCF because she would prefer the alternate SCF to $F$ under the deception $\alpha$, ruling out the deception as an equilibrium of a larger mechanism.
@@
We follow the standard timing used in implementation theory with the requirement that the signal realizations are observed sometime prior to when agents choose their strategies in the mechanism so that the informational content of the signal is factored into their strategies. [^6] Technically one can interpret an agent’s signal $\sigma_{i}$ about other agents as part of agent $i$’s private information, so if $\sigma_{i}$ has possible realizations in $\zeta_{i}=\{A_{i1},\dots,A_{ik}\}$ an implementing mechanism should ask each agent $i$ to specify a strategy for each $(\theta_{i},A_{ij})$ pair. Our setup is implicitly analyzing an equivalent formulation of this mechanism in which the uncertainty coming from the signals $\{\sigma_{i}\}$ is resolved first and asking whether $F$ is partially or fully implementable in each resulting subgame of the full mechanism. When the signals $\{\sigma_{i}\}_{i\in I}$ provide no information, the model is equivalent to the classical Bayesian implementation problem studied by Jackson (1991). The main question we address is whether and when an SCF $F$ that is not partially (resp. fully) implementable when $\sigma_{i}$ is noninformative for all $i$ can become partially (resp. fully) if one or more of the $\sigma_{i}$ is nontrivial. We are agnostic on where the signals $\{\sigma_{i}\}_{i\in I}$ come from. The nature of the question is one of comparative statics on the implementability of an SCF $F$ with respect to an increase in information one more agents $i$ has regarding $\Theta_{-i}$. [^7] Since agents initially do not have any information about $\Theta_{-i}$ aside from the common prior $\mu$, any nontrivial signal $\sigma_{i}$ represents an increase in agent $i$’s information with respect to the Blackwell order. In the remainder of this paper we will say that $F$ is partially implementable with respect to $\{\sigma_{i}\}_{i\in I}$ if truth-telling is an equilibrium in the direct mechanism in each realization of the signals $\{\sigma_{i}\}_{i\in I}$. Similarly define fully implementable with respect to $\{\sigma_{i}\}_{i\in I}.$
@@
Our setup makes several substantive assumptions. Other than requiring all $\sigma_{i}$ to have full support, we assume there are no restrictions over the structure of the signal agent $i$ can receive. Thus, agents do not have private information that can constrain the types of signals they can process. Additionally, we assume the signals are not costly in order to focus on the role of the additional information regarding other agents’ types. Finally, the structure of the signals $\{\sigma_{i}\}_{i\in I}$ is assumed to be common knowledge. The restriction that each $\sigma_{i}$ has full support is avoids complications to the type space $\Theta$ after agents receive their signals. In particular, any time an agent receives a signal that allows her to completely rule out any type profile, her set of possible types strictly enlarges. For example, if agents received a perfect signal about $\theta_{-i}$, each agent’s private type is the full state of the world $\theta$ in this model, which reduces the problem to the classical Nash implementation problem solved by Maskin (1999). [^8] In particular, incentive compatibility is a trivial constraint in the complete information setting. [^9] See Bergemann and Morris (2005) for a more detailed discussion of how changes to the type space affect implementation of SCFs.